OpenStudy (anonymous):

I will medal!! What is the solution to the equation -2(1+5a)=-6(a+3) ? a =

OpenStudy (welshfella):

first expand the 2 parentheses

OpenStudy (welshfella):

-2*2 -2*5a = -6*a - 6*3 can you continue?

OpenStudy (anonymous):

-14a=-24a

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

no you can only add/subtract like terms -4 -10a cannot be subtracted as one is a number and the other is a term in a

OpenStudy (welshfella):

-4 - 10a = -6a - 18 the next thing to do is to add 6a to both sides of the equation so as to get the 2 terms in a on the left side

OpenStudy (welshfella):

- 4 - 10a = - 6a - 18 +6a +6a

OpenStudy (welshfella):

what does that give us?

OpenStudy (welshfella):

-6a + 6a = ?

OpenStudy (anonymous):

0

OpenStudy (welshfella):

right so that gets rid of the -6a on the right side of our equation so we have - 4a -10a + 6a = -18

OpenStudy (welshfella):

sorry thats -4 - 10a + 6a = -18 -4 - 4a = -18 now to isolate -4a we have to add 4 to each side of the equation

OpenStudy (welshfella):

what will that give us? add 4 to left side and add 4 to the right side

OpenStudy (anonymous):

4a =-14

OpenStudy (welshfella):

almost there is a - in front of the 4a -4 + 4 - 4a = -18 + 4 -4a = -14

OpenStudy (welshfella):

now you divide both sides by -4 so that we get a only on left side. this way we get the solution. The value of a.

OpenStudy (welshfella):